OpenStudy (anonymous):

3 pizza and 4 sandwiches cost $34. 3 pizza and 7 sandwiches cost $41.50. How much does a pizza cost?

OpenStudy (anonymous):

3p+4s=34 3p+7s=41.5 Do I substitute?

OpenStudy (anonymous):

You can but this problem lends itself well to elimination if you multiply one of the equations by -1

OpenStudy (anonymous):

Okay thank you.

OpenStudy (anonymous):

I cqn also add them correct?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but how would I do that? Just regularily add? Or...

OpenStudy (anonymous):

Yes. you'd have add the like terms and the p's would cancel

OpenStudy (anonymous):

-3p-4s=-34 3p+7s=41.5 ______________ 3s = 7.5

OpenStudy (anonymous):

so...6P+11s=75.5

OpenStudy (anonymous):

You have to multiply one of them by -1 before adding them together.

OpenStudy (anonymous):

okay... in that case...

OpenStudy (anonymous):

-3p-7s=-41.5? Multiply all?

OpenStudy (anonymous):

-7.5

OpenStudy (anonymous):

Yes you multiply all of them

OpenStudy (anonymous):

You got -3s = -7.5?

OpenStudy (anonymous):

mhmm.

OpenStudy (anonymous):

39+30/12=34..

OpenStudy (anonymous):

and from here you can solve for s

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

So...

OpenStudy (anonymous):

17-5/2...right?

OpenStudy (anonymous):

i think...

OpenStudy (anonymous):

basically...

OpenStudy (anonymous):

wait wait wait..being a dumbcluck here.

OpenStudy (anonymous):

34/2 is...um...

OpenStudy (anonymous):

36/2!...

OpenStudy (anonymous):

... ><''

OpenStudy (anonymous):

The 5/2 is right.

OpenStudy (anonymous):

3p+7(5/2)=41.5

OpenStudy (anonymous):

We just substitute from there?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

3p+35/2=41.5 So what do I do then?

OpenStudy (anonymous):

solve for p. 3p = 41 - 35/2

OpenStudy (anonymous):

3p = 24

OpenStudy (anonymous):

okay..thanks.